# Offline Robotic World Model: Learning Robotic Policies without a Physics Simulator

### v1

@@ A.1.2 Reward functions
Symbol $w_{v_{xy}}$ $w_{\omega_{z}}$ $w_{v_{z}}$ $w_{\omega_{xy}}$ $w_{q_{\tau}}$
Value $1.0$ $0.5$ $-2.0$ $-0.05$ $-2.5e^{-5}$
Symbol $w_{\ddot{q}}$ $w_{\dot{a}}$ $w_{f_{a}}$ $w_{c}$ $w_{g}$
Value $-2.5e^{-7}$ $-0.01$ $0.5$ $-1.0$ $-5.0$
@@
where $c_{u}$ denotes the counts of the undesired knee contacts.
@@ A.3 Training Parameters
A.3.1 RWM-O
The training information of RWM-O is summarized in Table S8.
Table S8: RWM-O training parameters
@@ A.3.1 RWM-O
max iterations -- $2500$
learning rate -- $1e^{-4}$
weight decay -- $1e^{-5}$
batch size -- $1024$
ensemble size $B$ $5$
@@
transitions in training data -- $6M$
approximate training hours -- $1$
number of seeds -- $5$
@@ A.3.2 MOPO-PPO
imagination environments -- $4096$
imagination steps per iteration -- $100$
@@
max iterations -- $2500$
learning rate -- $0.001$
weight decay -- $0.0$
learning epochs -- $5$
mini-batches -- $4$
KL divergence target -- $0.01$
@@
entropy coefficient -- $0.005$
number of seeds -- $5$
A.4 Offline Policy Learning
A.4.1 Policy Deployment Learned with MOPO-PPO
The policy deployment learning with MOPO-PPO is visualized in Fig. S5.
Figure S5: MOPO-PPO policy deployment learned from offline Mixed dataset across various environments.
A.4.2 Real-World Data Collection
Although RWM-O can be trained on offline data from arbitrary sources, we develop an autonomous pipeline for collecting additional real-world data with minimal human intervention. The robot executes a velocity-tracking policy and records transitions as described in Sec. A.1.1. A visualization of the collection setup is shown in Fig. S6. To ensure safety, a rectangular boundary is defined relative to the odometry origin. Robot odometry is computed by fusing legged odometry and LiDAR data. At each sampling interval, a base velocity command (green arrow) is randomly drawn, and the resulting final pose (red sphere) is predicted via finite integration. If this predicted pose lies outside the safety boundary, a new command is sampled. After a fixed number of retries, a fallback command that drives the robot toward the origin is used instead. To mitigate odometry drift, the safety region can be re-centered around the robot’s current position using a joystick override. In total, we collect approximately one hour of data corresponding to 180K state transitions.
@@ A.4.2 Real-World Data Collection
A.4.3 Additional Offline Data Sources
RWM-O can achieve reasonable performance when trained exclusively on real-world data, largely due to the inherent distributional bias in the dataset. Since all trajectories are collected from an expert policy without failures such as falls or collisions, the world model only needs to generalize within a narrow, well-behaved region of the state-action space. MOPO-PPO capitalizes on this bias by penalizing transitions with high epistemic uncertainty, discouraging the policy from entering regions outside the dataset’s support. However, this mechanism alone is insufficient for fully stable policy learning. In particular, the absence of privileged signals, such as contact labels, limits the ability of RWM-O to predict physically grounded dynamics, especially during transitions involving contact events. These signals are critical for accurate reward computation, and their absence introduces noise in policy gradients, leading to instability during training. To mitigate this, additional offline data from simulation, with access to privileged information, can be incorporated to enhance the world model’s predictive capabilities. This improves the fidelity of reward estimation during imagination rollouts, resulting in more stable and robust policies that transfer reliably to hardware.

### v3

@@ A.1.2 Reward functions
Symbol Value Symbol Value Symbol Value Symbol Value
ANYmal D Unitree G1
$w_{v_{xy}}$ $1.0$ $w_{\omega_{z}}$ $0.5$ $w_{v_{xy}}$ $1.0$ $w_{\omega_{z}}$ $0.5$
$w_{v_{z}}$ 2.0-2.0 $w_{\omega_{xy}}$ 0.05-0.05 $w_{v_{z}}$ 2.0-2.0 $w_{\omega_{xy}}$ 0.05-0.05
$w_{q_{\tau}}$ $2.5e^{-5}$ $w_{\ddot{q}}$ $2.5e^{-7}$ $w_{q_{\tau}}$ $2.5e^{-5}$ $w_{\ddot{q}}$ $2.5e^{-7}$
$w_{\dot{a}}$ 0.01-0.01 $w_{f_{a}}$ $0.5$ $w_{\dot{a}}$ 0.05-0.05 $w_{f_{a}}$ $0.0$
$w_{c}$ 1.0-1.0 $w_{g}$ 5.0-5.0 $w_{c}$ 1.0-1.0 $w_{g}$ 5.0-5.0
$w_{f_{c}}$ $0.0$ $w_{q_{d}}$ $0.0$ $w_{f_{c}}$ $1.0$ $w_{q_{d}}$ 1.0-1.0
@@
where $c_{u}$ denotes the counts of the undesired contacts.
@@
Foot clearance
$r_{f_{c}}=w_{f_{c}}h_{f_{c}},$
where $h_{f_{c}}$ denotes the clearance height of the swing feet.
Joint deviation
$r_{q_{d}}=w_{q_{d}}\left\|q-q_{0}\right\|_{1},$
where $q_{0}$ denotes the default joint position.
@@ A.2.2 MOPO-PPO
A.2.3 Baselines
All MBRL baselines use the same architecture and hyperparameters for consistency and fair comparison as detailed in Table S9.
Table S9: Baseline architecture
Method MBPO MBPO-PPO MOPO MOPO-PPO
Model Architecture RWM RWM RWM-U RWM-U
Policy Optimization SAC PPO SAC PPO
@@ A.3 Training Parameters
A.3.1 RWM-U
The training information of RWM-U is summarized in Table S10.
Table S10: RWM-U training parameters
@@ A.3.1 RWM-U
max iterations - $2500$
learning rate - $1e^{-4}$
weight decay - $1e^{-5}$
batch size - $1024$
ensemble size $B$ 55
@@
transitions in training data - $6M$
approximate training hours - 11
number of seeds - 55
@@ A.3.2 MOPO-PPO
imagination environments - $4096$
imagination steps per iteration - $100$
@@
max iterations - $2500$
learning rate - $0.001$
weight decay - $0.0$
learning epochs - 55
mini-batches - 44
KL divergence target - $0.01$
@@
entropy coefficient - $0.005$
number of seeds - 55
A.4 Real-World Data Collection
Although RWM-U can be trained on offline data from arbitrary sources, we develop an autonomous pipeline for collecting additional real-world data with minimal human intervention. The robot executes a velocity-tracking policy and records transitions as described in Sec. A.1.1. A visualization of the collection setup is shown in Fig. S8. To ensure safety, a rectangular boundary is defined relative to the odometry origin. Robot odometry is computed by fusing legged odometry and LiDAR data. At each sampling interval, a base velocity command (green arrow) is randomly drawn, and the resulting final pose (red sphere) is predicted via finite integration. If this predicted pose lies outside the safety boundary, a new command is sampled. After a fixed number of retries, a fallback command that drives the robot toward the origin is used instead. To mitigate odometry drift, the safety region can be re-centered around the robot’s current position using a joystick override. In total, we collect approximately one hour of data corresponding to 200K state transitions.
@@ A.4 Real-World Data Collection
A.5 Distinct Challenges
A large body of offline RL research is built upon off-policy, value-function-based methods such as CQL (Kumar et al., 2020), IQL (Kostrikov et al., 2021), and COMBO (Yu et al., 2021). While these approaches achieve strong results on standard offline RL benchmarks, they do not translate directly to high-dimensional real-world locomotion. In contemporary legged robotics, there is substantial empirical evidence that on-policy policy-gradient methods, particularly PPO (Schulman et al., 2017), provide significantly more robust and reliable performance. Although the precise reasons for this difference remain an active research topic, several observations have emerged from practice. Modern locomotion systems rely on massively parallel simulation, where wall-clock training speed rather than sample efficiency dominates the design. PPO’s simple update structure and stable optimization behavior make it effective in this regime, whereas off-policy critics such as SAC (Haarnoja et al., 2018) must estimate and update complex value networks that are more difficult to stabilize when scaled to thousands of environments. Off-policy value-based methods also tend to struggle with the distributional shift characteristic of locomotion control. Real-world proprioceptive trajectories are highly correlated, and critic-based objectives are prone to overfitting narrow regions of the state–action space induced by the most recent data distribution. This can lead to brittle policies that generalize poorly and do not transfer reliably to hardware. Existing investigations further highlight the difficulty of tuning SAC-based methods to match PPO’s stability in large-scale locomotion training (Raffin, 2025). Despite ongoing research, there is currently no robust recipe for deploying SAC-style approaches in this domain.
These considerations directly constrain the design of offline RL methods capable of operating on physical robots. Offline algorithms built on off-policy value functions cannot be readily applied to locomotion because the resulting policies are not reliably deployable. Moreover, many value-based offline RL approaches are not model-based and therefore cannot leverage synthetic rollouts—an essential capability when no further real-world data can be collected. For locomotion tasks involving long-horizon credit assignment and limited dataset coverage, the ability to generate additional imagined trajectories through a learned dynamics model is crucial. Consequently, this work focuses on model-based and on-policy formulations, which better align with the requirements of fully offline learning in real high-dimensional robotic systems.
Within this context, RWM-U addresses several distinctive challenges inherent to fully offline model-based reinforcement learning. Prior offline MBRL approaches, such as MOPO (Yu et al., 2020), have largely been evaluated in simplified simulated settings and typically employ single-step feedforward ensembles combined with short-horizon value-based optimization. In contrast, our method builds on the autoregressive world model introduced in RWM, enabling long-horizon dynamics learning suitable for locomotion control. A central aspect of our approach is the ability to quantify and propagate epistemic uncertainty over extended model-generated trajectories. Ensemble disagreement naturally reflects compounding prediction error in regions with limited data support, as illustrated in Fig. 3 (right), and provides a principled mechanism for avoiding unreliable parts of the state–action space.
We further develop MOPO-PPO, which adapts MOPO’s uncertainty-penalized training framework to on-policy PPO. This adaptation is nontrivial: unlike MOPO’s short one-step rollouts with SAC, our system must incorporate uncertainty-aware penalties over autoregressive rollouts spanning up to 100 steps. Maintaining stability in the presence of compounding model uncertainty under a strict fully offline constraint introduces algorithmic challenges distinct from those faced in online or partially interactive settings. Finally, we demonstrate that this uncertainty-aware long-horizon offline MBRL pipeline can be deployed on a physical quadruped and humanoid, where policies are trained entirely offline from a fixed dataset. To our knowledge, this represents the first demonstration of uncertainty-penalized long-horizon offline MBRL operating reliably on real robotic hardware.
A.6 Online vs. Offline Reinforcement Learning
Reinforcement learning can be conducted under several interaction regimes, and the distinction between online and offline settings has important implications for both algorithm design and achievable performance. In online RL, the learning system is allowed to interact with the environment throughout training. After each policy or model update, the controller can return to the agent to collect additional trajectories, which serve as fresh supervision. This iterative feedback loop enables the system to continuously correct model errors, reduce domain mismatch between predicted and real dynamics, and expand the coverage of the experienced state–action distribution as the policy evolves. Recent robotics works that demonstrate strong real-world learning performance operate within this paradigm (Yang et al., 2020; Smith et al., 2022; Levy et al., 2024). Although some training phases in these works occur offline, each method continually acquires new on-robot data, allowing the model and policy to adapt whenever they become unreliable.
In contrast, the fully offline RL setting considered in this work assumes that no additional environment interaction is permitted once the dataset is collected (Levine et al., 2020). All components of the system—dynamics modeling, long-horizon rollouts, and policy optimization—must operate exclusively on a fixed dataset. This imposes fundamentally different constraints. The dataset may contain only a limited number of behaviors, be biased toward safe or suboptimal trajectories, and lack critical transitions such as failures or recovery maneuvers that are naturally observed during online exploration. Without the ability to gather new samples, the learning system cannot correct model inaccuracies or distributional biases by expanding its coverage of the state–action space. As a result, distribution shift is substantially amplified, model prediction errors accumulate over long autoregressive rollouts, and the policy lacks any external feedback when entering underrepresented regions of the data (Levine et al., 2020; Prudencio et al., 2023).
These differences lead to distinct algorithmic requirements. Whereas online RL can rely on continual data collection to stabilize training, offline RL must depend solely on mechanisms internal to the model and policy optimization process. In particular, uncertainty estimation becomes essential for detecting underexplored states and preventing the policy from drifting into regions where the model is unreliable. The uncertainty-aware design of RWM-U and MOPO-PPO is therefore motivated directly by the challenges inherent to fully offline settings rather than by those faced in online or iterative data-collection regimes.
The contrast between online and offline RL is thus not a minor technical distinction but a defining aspect of the problem formulation. Online methods leverage real-world feedback to mitigate compounding errors and bridge domain gaps, whereas fully offline methods must succeed without such corrective interaction. This work focuses on the latter regime and investigates the design principles required to make fully offline model-based reinforcement learning viable for real robotic systems.